OpenStudy (anonymous):

HELP PLZ WILL MEDAL, Solve the given number problem. The Sum of three consecutive even integers is -72 what is the least of these numbers? -26 -22 -30 24

OpenStudy (anonymous):

the options are -26 -22 -30 +24

OpenStudy (anonymous):

@marigirl @marigirl @marigirl

OpenStudy (anonymous):

@ganeshie8 @ganeshie8 @ganeshie8

OpenStudy (anonymous):

Thanks mar...i really need help

OpenStudy (marigirl):

sum of three consecutive EVEN integers x x+2 x+4

OpenStudy (anonymous):

\[x+(x+2(x+4))\] ...So an equation like this i guess

OpenStudy (marigirl):

yes and that will equal -72

OpenStudy (anonymous):

divide -72 by three to find the middle number

OpenStudy (anonymous):

well im not sure....i dont really understnad what it means by least of these numbers...what nums should i plug in for x

OpenStudy (anonymous):

72/3 =24

OpenStudy (anonymous):

-24, and what is one number lower?

OpenStudy (anonymous):

thus making it \[x+(24+2(x+4))\] ...-22 i guess

OpenStudy (anonymous):

-23

OpenStudy (anonymous):

oh oops ok...so what now?

OpenStudy (marigirl):

NO i got \[x+x+2+x+4=-72\] \[3x+6=-72\] \[3x=-78\] \[x=-26\] the least means that it should be the lowest

OpenStudy (anonymous):

I just saw even integers sorry -22,-24-26=-72 are the three Which is the least

OpenStudy (anonymous):

-22

OpenStudy (anonymous):

since they are negative -26 would be the lowest

OpenStudy (anonymous):

lowest = closest to zero i think....

OpenStudy (marigirl):

no lowest means closer to negative infinity

OpenStudy (anonymous):

ok....but what about -30....thats farthest and its a negative

OpenStudy (anonymous):

When they are positive numbers the that would be correct. Keep counting backwards and which is the furthest from zero?

OpenStudy (marigirl):

x+x+2+x+4=−72 3x+6=−72 3x=−78 x=−26 ur numbers are -26 -24 -22 all this will add to -72 the least means that it should be the lowest

OpenStudy (anonymous):

It asked for 3 consecutive numbers that added up to -72 which numbers are those? those 3 will be the only ones to consider

OpenStudy (anonymous):

AHHHHHHH...ok now i get this...jesus christ im thick as maple syrup

OpenStudy (anonymous):

THx guys for the help....

OpenStudy (anonymous):

OMF it was right thanks guys...big help...

OpenStudy (anonymous):

:) :D